OpenStudy (anonymous):

A square pyramid has a slant height of 17 inches and a height of 15 inches. What is its surface area if the square base has sides of length 16 inches? Answer 224 in2 800 in2 864 in2 480 in2

OpenStudy (anonymous):

why cant the pic a shape and stay with it

OpenStudy (anonymous):

what do you mean sunset?

OpenStudy (anonymous):

i lurn how to do one and them they change the question and i get lost again

OpenStudy (anonymous):

they mix it up to much

OpenStudy (anonymous):

aww, actually it's not that difficult, just some practice and visualization.

OpenStudy (anonymous):

its hard for me to get the formula do for all the shape i can get one and then they change it its confusing for me

OpenStudy (anonymous):

For this the answer is \( \boxed {2} \)

OpenStudy (anonymous):

will give me the formula for this one to please

OpenStudy (anonymous):

\( \frac 12 \times \) Perimeter of the base \( \times \) slant height + Area of the base.

OpenStudy (anonymous):

thank u again im sorry for all the questions

OpenStudy (anonymous):

Glad to help and not need to apologize :) If you have more just shoot :)

OpenStudy (anonymous):

couldnt you just work the area of the base as a square, then the triangle are x4 add them up, but i guess the formula is much quicker